### v1

@@ Abstract
Large language models (LLMs) have achieved significant performance progress in various natural language processing applications. However, LLMs still struggle to meet the strict requirements for accuracy and reliability in the medical field and face many challenges in clinical applications. Existing clinical diagnostic evaluation benchmarks for evaluating medical agents powered by LLMs have severe limitations. Firstly, most existing medical evaluation benchmarks face the risk of data leakage or contamination. Secondly, existing benchmarks often neglect the characteristics of multiple departments and specializations in modern medical practice. Thirdly, existing evaluation methods are limited to multiple-choice questions, which do not align with the real-world diagnostic scenarios. Lastly, existing evaluation methods lack comprehensive evaluations of end-to-end real clinical scenarios. These limitations in benchmarks in turn obstruct advancements of LLMs and agents for medicine. To address these limitations, we introduce ClinicalLab, a comprehensive clinical diagnosis agent alignment suite. ClinicalLab includes ClinicalBench, an end-to-end multi-departmental clinical diagnostic evaluation benchmark for evaluating medical agents and LLMs. ClinicalBench is based on real cases that cover 24 departments and 150 diseases. We ensure that ClinicalBench does not have data leakage. ClinicalLab also includes four novel metrics (ClinicalMetrics) for evaluating the effectiveness of LLMs in clinical diagnostic tasks. We evaluate 17 general and medical-domain LLMs and find that their performance varies significantly across different departments. Based on these findings, in ClinicalLab, we propose ClinicalAgent, an end-to-end clinical agent that aligns with real-world clinical diagnostic practices. We systematically investigate the performance and applicable scenarios of variants of ClinicalAgent on ClinicalBench. Our findings demonstrate the importance of aligning with modern medical practices in designing medical agents. The code and dataset are publicly available at https://github.com/WeixiangYAN/ClinicalLab.
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated exceptional capabilities across a wide variety of natural language processing tasks [12, 37, 9, 30, 33]. Recent studies attempt to apply LLMs to the medical field [29], aiming to improve the efficiency of healthcare systems through early disease diagnosis and timely health management, thereby reducing the workload of medical professionals. However, the high sensitivity of the medical field and the strict requirements for accuracy and reliability in clinical decision-making pose great challenges to LLMs. Recent studies find that LLMs are prone to producing hallucinations in medical diagnoses [28, 4], and such erroneous diagnoses could harm the physical and psychological health of patients and also potentially lead to serious medical negligence. How to effectively, reliably, and comprehensively evaluate the true capabilities of LLMs and the accuracy of their diagnostic results, as well as reveal their limitations to avoid potential misdiagnosis, remains an unresolved problem.
Recent studies find that existing benchmarks cannot effectively evaluate the medical capabilities of LLMs [43, 36]. Firstly, existing benchmarks are often based on data collected from online consultation platforms or medical textbooks [34], which could easily be included in the training data of LLMs, that is, leading to data leakage or contamination and thus biasing the performance evaluation of LLMs. Secondly, the departmental setup in modern medicine is designed to address the complex medical needs of different structures and functions of human organs. The specific skills and treatment methods vary significantly across different departments. However, existing evaluation benchmarks overlook the characteristics of multi-departmental and highly specialized nature of modern medicine, hence they are insufficient in capturing performance differences across departments. Thirdly, existing evaluation methods typically confine themselves to multiple-choice questions [14, 24, 44], which does not align with real-world clinical diagnostic scenarios. In actual medical environments, patients seek medical services because they are uncertain about their health conditions, rather than knowing the possible disease options and then seeking a doctor’s judgment. Last but not least, there is currently no evaluation method that can comprehensively and reliably evaluate the end-to-end practicality of LLMs in the entire clinical diagnosis process, which starts from the moment a patient enters the clinic and ends when the patient is discharged. This issue will in turn limit the design and evaluation of practical medical agents powered by LLMs and harm exploitation of the full potential of LLMs.
@@
Inspired by the findings, we propose ClinicalAgent, a novel clinical diagnostic agent that automatically assigns departments based on patient complaints and organizes joint consultations with the top-K experts in the assigned department to provide unified diagnostic results. Figure 1 depicts the workflow of ClinicalAgent. Experiments show that ClinicalAgent outperforms top-performing LLMs in clinical diagnostic performance, demonstrating the importance of aligning with modern medical practice for advancing agentic medical systems. We release ClinicalLab, a comprehensive clinical diagnosis agent alignment suite, including ClinicalBench, ClinicalMetrics, and ClinicalAgent, to promote development of clinical diagnostic agents.
Our contributions can be summarized as follows:
  • ClinicalBench: We establish the first real-case-based, data-leakage-free, end-to-end multi-departmental benchmark for evaluating the clinical diagnostic capabilities of LLMs. This benchmark covers most departments (totalling 24) and most extensive varieties of diseases (totalling 150).
  • ClinicalMetrics: We introduce four novel metrics to evaluate the effectiveness of LLMs in clinical diagnostics, including Department Win Rate, Department Instruction Following Rate, Comprehensive Diagnostic Accuracy, and Acceptability.
@@ 2.1 Existing Medical Benchmarks
The Chinese Medical Benchmark [34] consists of CMB-Exam and CMB-Clin, with data collected from various public examination databases and multiple-choice questions from textbooks. MedBench [5] also includes multiple-choice questions from the Chinese mainland medical licensing examination and 1025 QA pairs based on electronic medical records. MMedBench [26] is a multilingual medical evaluation benchmark covering six languages, with data sourced from medical textbooks and open-source medical websites in different languages. Different from our method of collecting real cases from different departments, MMedBench and CMExam [20] utilize GPT-4 to categorize each question according to a pre-defined list of departments. However, according to the experimental results shown in Figure 4, GPT-4 exhibits significant errors in departmental classification. MedQA [15] consists of multiple-choice questions collected from professional medical board examinations. Its USMLE-style English subset offers four possible answer options for each question and is widely used to evaluate the performance of LLMs in the medical domain. PubMedQA [16] is a biomedical QA dataset collected from the PubMed website, where questions need to be answered with Yes/No/Maybe. MMLU [12] is a benchmark covering 57 tasks across multiple domains, recent work [28] extracts 6 medicine-related tasks from MMLU to evaluate the medical capabilities of LLMs. Furthermore, HealthSearchQA [28] is a dataset based on medical conditions and related symptoms frequently searched by users on search engines, aiming to evaluate a model’s ability to handle real-world user’s concerns about their health.
Table 1 compares statistics of existing benchmarks and datasets for evaluating LLMs in the medical field and our ClinicalBench, including the number of samples, task types, language coverage, and data sources. In summary, the main shortcomings of existing evaluation benchmarks include: (1) lack of comprehensive and evenly distributed departmental coverage to prevent evaluation bias; (2) data sources often come from easily accessible online consultation platforms, medical textbooks, and professional examinations, which poses high risks of data leakage; (3) existing benchmarks primarily test medical knowledge through multiple-choice questions, which differ significantly from real-world diagnostic scenarios.
@@ 2.2 Existing Agents for Medical Applications
Recent works attempt to solve medical and healthcare issues through paradigms of dividing tasks and collaboration among multiple agents. MedAgent [31] is the first multi-agent framework in the medical field that improves the accuracy of solving medical multiple-choice questions by allowing the same LLM to play different roles in multi-round collaborative dialogues. However, the design of MedAgent still relies on the multiple-choice question format, which differs significantly from the diagnostic process in the real world. On the other hand, Williams et al. [35] effectively simulate real-world human behavior patterns through agents played by ChatGPT to address the challenge of incorporating human behavioral factors into epidemic models. MDAgents [17] is a multi-agent framework that employs an adaptive decision-making mechanism, addressing medical multiple-choice questions through multiple stages of checking problem complexity, dynamically recruiting experts, reasoning, and decision-making. Meanwhile, Agent Hospital [18] improves diagnostic accuracy by simulating nearly all medical processes within a hospital and trains doctor agents through doctor-patient interaction simulations. However, due to the lack of a comprehensive dataset covering the entire medical process, currently its effectiveness is only validated on the MedQA multiple-choice dataset. In summary, existing medical agents suffer from severe limitations and constraints in designs and evaluations due to the lack of benchmarks and datasets based on real diagnostic processes. Therefore, developing an evaluation benchmark based on real medical diagnostic processes is critical. Furthermore, most previously proposed medical agents perform clinical diagnosis by using prompts to make the same model play different roles. In contrast, our ClinicalAgent is driven by the evaluation results and analysis of prior models on ClinicalBench. ClinicalAgent uses different models to play different roles and implements dynamic department scheduling and doctor allocation strategies for clinical diagnosis, ensuring a strong alignment with real hospital settings.
@@ 3.1 Data Sources & Licenses
The data samples used in the ClinicalBench benchmark are sourced from real clinical medical records of officially certified Grade 3A hospitals in China [^1] Grade 3A hospitals are the highest level hospitals in China’s “three-grade, six-class” classification system.. The collection of this data strictly adheres to the principles of patient privacy protection. No information related to the hospitals is disclosed. As detailed in Data Processing & Quality, to protect patient privacy, any personally identifiable information (PII) of patients, treatment regions, or other sensitive information has been manually identified and removed by the team of doctors. All data is obtained legally and ethically, and has been reviewed and approved by the Ethics Committees of the relevant hospitals, ensuring that research activities on these data comply with ethical and legal obligations.
We are committed to responsible data management and strictly follow relevant laws and regulations involving the collection, use, and distribution of protected health information. To ensure the legal and regulated use of the dataset, we have formulated the ClinicalBench Usage and Data Distribution License Agreement, which can be found in the Github. This agreement strictly requires all users to use the data solely for research purposes and to adhere to strict regulations protecting patient privacy, prohibiting any form of personal information tracking or identification. Through these measures, we ensure the legality and ethics of data acquisition and use while supporting research that may promote the development of LLMs in clinical diagnostics.
@@ 3.3 Data Statistics
ClinicalBench is a fine-grained evaluation benchmark specifically designed for multi-departmental clinical diagnosis, covering 24 departments such as pediatrics, orthopedics, and neurosurgery. Figure 4 presents detailed information about the various departments covered by ClinicalBench. It involves 150 different diseases, each comprising 10 specific cases, totaling 1500 samples, with an average of about 1000 tokens per case [^2] All samples are originally in Chinese and translated into English using GPT-4 for reading and usage.. Table 1 provides relevant information about ClinicalBench. To the best of our knowledge, ClinicalBench is the most comprehensive clinical diagnostics evaluation benchmark to date, covering the widest range of departments and diseases.
@@
Figure 4: Ranking of different LLMs across departments, with the x-axis representing department abbreviations (abbreviations correspond to Figure 4) and the y-axis representing the models.
Each case in ClinicalBench contains detailed clinical data, such as the patient’s gender, age [^3] Gender and age are crucial factors in clinical diagnosis, and with PII and other sensitive information removed, they alone are typically not considered PII. Therefore, we decide to retain these two information., chief complaint, medical history, and physical examination. Additionally, the cases include various medical imaging reports, such as X-rays, computed tomography (CT) scans, magnetic resonance imaging (MRI), and ultrasound examinations, as well as biochemical, immunological, microbiological, and pathological laboratory examination results from biological samples such as blood, urine, and cerebrospinal fluid. The preliminary diagnosis, diagnostic basis, differential diagnosis, and treatment plans provided by doctors from each department collectively support a range of end-to-end clinical decision-making processes. A complete data example is provided in Table 5-9.
@@ 3.4 Task Overview
3.4.1 Department Guide (Multi-Choice QA)
When patients first enter the hospital, guiding them to the correct department for further consultation, examination and treatment is crucial for providing targeted medical services. This task requires the model to choose the most reasonable department from a given set of 24 departments based on a chief complaint $CC$, i.e., $D=\text{LLM}_{\textit{dept\_guide}}(CC)$, where $D$ denotes the selected department. This task evaluates the decision-making accuracy, and instruction-following capability of LLMs in real-world medical scenarios.
3.4.2 Clinical Diagnosis (Generative QA & Multi-Choice QA)
The Clinical Diagnosis task evaluates the LLM’s clinical diagnostic capabilities in real cases and its analysis and interpretation abilities. Clinical Diagnosis includes 6 sub-tasks, each designed according to the standard diagnostic process recommended by the Chinese official record [22]. Specifically, each case $K$ includes chief complaint $CC$, medical history $H$, physical examination $V$, laboratory examination results $A$, and medical imaging reports $R$. The 6 sub-tasks are defined as follows.
Preliminary Diagnosis (PD): The model generates a list of possible diseases $P$ based on $CC$, $H$, $V$, and $A$, i.e., $P=\text{LLM}_{\textit{PD}}(CC,H,V,A)$. This task primarily evaluates the model’s information synthesis and reasoning abilities.
Diagnostic Basis (DB): The model needs to provide supportive medical evidence $B$ for each possible disease in $P$, i.e., $B=\text{LLM}_{\textit{DB}}(P,K)$. This task mainly evaluates the model’s evidence extraction ability and helps mitigate hallucinations.
Differential Diagnosis (DD): After considering $P$, the model should perform a differential analysis to exclude other diseases $D$ that have different causes but similar manifestations, i.e., $D=\text{LLM}_{\textit{DD}}(P,B,K)$. This task primarily evaluates the model’s abilities in analytical comparison and decision-making.
Final Diagnosis (FD): Integrating all information, the model needs to determine the final diagnosis $F$, i.e., $F=\text{LLM}_{\textit{FD}}(P,D,B,K)$. This task mainly evaluates the model’s comprehensive judgment ability and the accuracy of clinical diagnosis.
Principle of Treatment (PT): The model determines the treatment principles and guidelines $G$ for $F$, i.e., $G=\text{LLM}_{\textit{PT}}(F)$. This task primarily evaluates the model’s memorization of medical standard operating procedures and knowledge points.
Treatment Plan (TP): The model formulates specific treatment steps and plans $T$ for the given case $K$, including medication, surgical intervention, and physical therapy, i.e., $T=\text{LLM}_{\textit{TP}}(F,G,K)$. This task mainly evaluates the model’s abilities in knowledge application and strategic planning.

### v2

@@ Abstract
Large language models (LLMs) have achieved significant performance progress in various natural language processing applications. However, LLMs still struggle to meet the strict requirements for accuracy and reliability in the medical field and face many challenges in clinical applications. Existing clinical diagnostic evaluation benchmarks for evaluating medical agents powered by LLMs have severe limitations. Firstly, most existing medical evaluation benchmarks face the risk of data leakage or contamination. Secondly, existing benchmarks often neglect the characteristics of multiple departments and specializations in modern medical practice. Thirdly, existing evaluation methods are limited to multiple-choice questions, which do not align with the real-world diagnostic scenarios and are not robust. Lastly, existing evaluation methods lack comprehensive evaluations of end-to-end real clinical scenarios. These limitations in benchmarks in turn obstruct advancements of LLMs and agents for medicine. To address these limitations, we introduce ClinicalLab, a comprehensive clinical diagnosis agent alignment suite. ClinicalLab includes ClinicalBench, an end-to-end multi-departmental clinical diagnostic evaluation benchmark for evaluating medical agents and LLMs. ClinicalBench is based on real cases that cover 24 departments and 150 diseases. We ensure that ClinicalBench does not have data leakage. ClinicalLab also includes four novel metrics (ClinicalMetrics) for evaluating the effectiveness of LLMs in clinical diagnostic tasks. We evaluate 17 general and medical-domain LLMs and find that their performance varies significantly across different departments. Based on these findings, in ClinicalLab, we propose ClinicalAgent, an end-to-end clinical agent that aligns with real-world clinical diagnostic practices. We systematically investigate the performance and applicable scenarios of variants of ClinicalAgent on ClinicalBench. Our findings demonstrate the importance of aligning with modern medical practices in designing medical agents. The code and dataset are publicly available at https://github.com/WeixiangYAN/ClinicalLab.
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated exceptional capabilities across a wide variety of natural language processing tasks [12, 9, 30]. Recent studies attempt to apply LLMs to the medical field [29], aiming to improve the efficiency of healthcare systems through early disease diagnosis and timely health management, thereby reducing the workload of medical professionals. However, the high sensitivity of the medical field and the strict requirements for accuracy and reliability in clinical decision-making pose great challenges to LLMs. Recent studies find that LLMs are prone to producing hallucinations in medical diagnoses [28, 4], and such erroneous diagnoses could harm the physical and psychological health of patients and also potentially lead to serious medical negligence. How to effectively, reliably, and comprehensively evaluate the true capabilities of LLMs and the accuracy of their diagnostic results, as well as reveal their limitations to avoid potential misdiagnosis, remains an unresolved problem.
Recent studies find that existing benchmarks cannot effectively evaluate the medical capabilities of LLMs [42, 35]. (1) Firstly, existing benchmarks are often based on data collected from online consultation platforms or medical textbooks [33], which could easily be included in the training data of LLMs, that is, leading to data leakage or contamination and thus biasing the performance evaluation of LLMs. (2) Secondly, the departmental setup in modern medicine is designed to address the complex medical needs of different structures and functions of human organs. The specific skills and treatment methods vary significantly across different departments. However, existing evaluation benchmarks overlook the characteristics of multi-departmental and highly specialized nature of modern medicine, hence they are insufficient in capturing performance differences across departments. (3) Thirdly, existing evaluation methods typically confine themselves to multiple-choice questions [14, 24, 43], which does not align with real-world clinical diagnostic scenarios. In actual medical environments, patients seek medical services because they are uncertain about their health conditions, rather than knowing the possible disease options and then seeking a doctor’s judgment. Besides, recent studies [40] have demonstrated that using multiple-choice questions to evaluate LLMs is not robust and introduces significant biases. (4) Last but not least, there is currently no evaluation method that can comprehensively and reliably evaluate the end-to-end practicality of LLMs in the entire clinical diagnosis process, which starts from the moment a patient enters the clinic and ends when the patient is discharged. This issue will, in turn, limit the design and evaluation of practical medical agents powered by LLMs and harm the exploitation of the full potential of LLMs.
@@
Inspired by the findings, we propose ClinicalAgent, a novel clinical diagnostic agent that dynamically allocates the $\mathcal{K}$ most relevant departments and assigns $\mathcal{N}$ clinicians from each department for a collaborative consultation based on the patient’s chief complaint. Leveraging flexible department scheduling and clinician allocation strategies, ClinicalAgent delivers a unified diagnostic result. Figure 1 depicts the workflow of ClinicalAgent. Experiments show that ClinicalAgent outperforms top-performing LLMs in clinical diagnostic performance, demonstrating the importance of aligning with modern medical practice for advancing agentic medical systems. We release ClinicalLab, a comprehensive clinical diagnosis agent alignment suite, including ClinicalBench, ClinicalMetrics, and ClinicalAgent, to promote development of clinical diagnostic agents. Our contributions can be summarized as follows:
  • ClinicalBench: We establish the first real-case-based, data-leakage-free, end-to-end multi-departmental benchmark for evaluating the clinical diagnostic capabilities of LLMs. This benchmark covers most departments (totalling 24) and most varieties of diseases (totalling 150).
  • ClinicalMetrics: We introduce four novel metrics to evaluate the practicality and effectiveness of the results generated by LLMs in real clinical scenarios.
@@ 2.1 Existing Medical Benchmarks
The Chinese Medical Benchmark [33] consists of CMB-Exam and CMB-Clin, with data collected from various public examination databases and multiple-choice questions from textbooks. MedBench [5] also includes multiple-choice questions from the Chinese mainland medical licensing examination and 1,025 QA pairs based on electronic medical records. MMedBench [26] is a multilingual medical evaluation benchmark covering six languages, with data sourced from medical textbooks and open-source medical websites in different languages. Different from our method of collecting real cases from different departments, MMedBench and CMExam [20] utilize GPT-4 to categorize each question according to a pre-defined list of departments. However, according to the experimental results shown in Table 4, GPT-4 exhibits significant errors in departmental classification. MedQA [15] consists of multiple-choice questions collected from professional medical board examinations. Its USMLE-style English subset offers four possible answer options for each question and is widely used to evaluate the performance of LLMs in the medical domain. PubMedQA [16] is a biomedical QA dataset collected from the PubMed website, where questions need to be answered with Yes/No/Maybe. MMLU [12] is a benchmark covering 57 tasks across multiple domains, recent work [28] extracts 6 medicine-related tasks from MMLU to evaluate the medical capabilities of LLMs. Furthermore, HealthSearchQA [28] is a dataset based on medical conditions and related symptoms frequently searched by users on search engines, aiming to evaluate a model’s ability to handle real-world user’s concerns about their health.
Table 1 compares statistics of existing benchmarks and datasets for evaluating LLMs in the medical field and our ClinicalBench, including the number of samples, task types, language coverage, and data sources. In summary, the main shortcomings of existing evaluation benchmarks include: (1) lack of end-to-end and evenly distributed departmental coverage to prevent evaluation bias; (2) data sources often come from easily accessible online consultation platforms, medical textbooks, and professional examinations, which poses high risks of data leakage; (3) existing benchmarks primarily evaluate medical knowledge through multiple-choice questions, which not only differ significantly from real-world diagnostic scenarios but also lack robustness.
@@ 2.2 Existing Agents for Medical Applications
Recent works attempt to solve medical and healthcare issues through paradigms of dividing tasks and collaboration among multiple agents. MedAgent [31] is the first multi-agent framework in the medical field that improves the accuracy of solving medical multiple-choice questions by allowing the same LLM to play different roles in multi-round collaborative dialogues. However, the design of MedAgent still relies on the multiple-choice question format, which differs significantly from the diagnostic process in the real world. On the other hand, williams2023epidemic [34] effectively simulate real-world human behavior patterns through agents played by ChatGPT to address the challenge of incorporating human behavioral factors into epidemic models. MDAgents [17] is a multi-agent framework that employs an adaptive decision-making mechanism, addressing medical multiple-choice questions through multiple stages of checking problem complexity, dynamically recruiting experts, reasoning, and decision-making. Meanwhile, Agent Hospital [18] improves diagnostic accuracy by simulating nearly all medical processes within a hospital and trains doctor agents through doctor-patient interaction simulations. However, due to the lack of a comprehensive dataset covering the entire medical process, currently its effectiveness is only validated on the MedQA multiple-choice dataset. In summary, existing medical agents suffer from severe limitations and constraints in designs and evaluations due to the lack of benchmarks and datasets based on real diagnostic processes. Therefore, developing an evaluation benchmark based on real medical diagnostic processes is critical. Furthermore, most previously proposed medical agents perform clinical diagnosis by using prompts to make the same model play different roles. In contrast, our ClinicalAgent is driven by the evaluation results and analysis of prior models on ClinicalBench. ClinicalAgent uses different models to play different roles and implements dynamic department scheduling and doctor allocation strategies for clinical diagnosis, ensuring a strong alignment with real hospital settings.
@@ 3.1 Data Sources & Licenses
The data samples used in the ClinicalBench benchmark are sourced from real clinical medical records of officially certified Grade 3A hospitals in China [^1] Grade 3A hospitals are the highest level hospitals in China’s “three-grade, six-class” classification system.. The collection of this data strictly adheres to the principles of patient privacy protection. No information related to the hospitals is disclosed. As detailed in Data Processing & Quality, to protect patient privacy, any personally identifiable information (PII) of patients, treatment regions, or other sensitive information has been manually identified and removed by the team of doctors. All data is obtained legally and ethically, and has been reviewed and approved by the Ethics Committees of the relevant hospitals, ensuring that research activities on these data comply with ethical and legal obligations. In Appendix 8, we discuss the certification documents issued by medical institutions and notary offices.
We are committed to responsible data management and strictly follow relevant laws and regulations involving the collection, use, and distribution of protected health information. To ensure the legal and regulated use of the dataset, we have formulated the ClinicalBench Usage and Data Distribution License Agreement, which can be found in the supplementary materials. This agreement strictly requires all users to use the data solely for research purposes and to adhere to strict regulations protecting patient privacy, prohibiting any form of personal information tracking or identification. Through these measures, we ensure the legality and ethics of data acquisition and use while supporting research that may promote the development of LLMs in clinical diagnostics. Appendix A.1 provides a detailed explanation of how we securely distribute the dataset.
@@ 3.3 Data Statistics
ClinicalBench is a fine-grained evaluation benchmark based on chain-of-thought reasoning, specifically designed for multi-departmental clinical diagnosis, covering 24 departments such as pediatrics, orthopedics, and neurosurgery. Figure 4 presents detailed information about the various departments covered by ClinicalBench. It involves 150 different diseases, each comprising 10 specific cases, totaling 1500 samples, with an average of about 1000 tokens per case [^2] All samples are originally in Chinese and translated into English using GPT-4 for reading and usage.. Table 1 provides relevant information about ClinicalBench. To the best of our knowledge, ClinicalBench is the most comprehensive clinical diagnostics evaluation benchmark to date, covering the widest range of departments and diseases.
@@
Figure 4: Ranking of different LLMs across departments, with the x-axis representing department abbreviations (abbreviations correspond to Figure 4) and the y-axis representing the models. The ranking of each department is determined based on the Avg. Score metric described in Table 2.
Each case $\mathcal{E}$ in ClinicalBench contains detailed clinical data, such as the patient’s gender, age [^3] Gender and age are crucial factors in clinical diagnosis, and with PII and other sensitive information removed, they alone are typically not considered PII. Therefore, we decide to retain these two information., chief complaint $\mathcal{CC}$, medical history $\mathcal{MH}$, physical examination $\mathcal{PE}$. Additionally, the cases include various medical imaging reports $\mathcal{IR}$, such as X-rays, computed tomography (CT) scans, magnetic resonance imaging (MRI), and ultrasound examinations, as well as biochemical, immunological, microbiological, and pathological laboratory examination reports $\mathcal{LR}$ from biological samples such as blood, urine, and cerebrospinal fluid. The preliminary diagnosis, diagnostic basis, differential diagnosis, and treatment plans provided by doctors from each department collectively support a range of end-to-end clinical decision-making processes. A complete data example is provided in Table 5-9.
@@ 3.4 Task Overview
3.4.1 Department Guide (Multi-Choice QA with 24 options)
When patients first enter the hospital, guiding them to the correct department for further consultation, examination and treatment is crucial for providing targeted medical services. The Department Guide (DG) task requires the model to choose the most reasonable department from a given set of 24 departments based on a $\mathcal{CC}$, i.e., $\mathcal{DP}=\text{DG}(\mathcal{CC})$, where $\mathcal{DP}$ denotes the selected department. This task evaluates the decision-making accuracy, and instruction-following capability of LLMs in real-world medical scenarios.
3.4.2 Clinical Diagnosis (Generative QA)
The Clinical Diagnosis task evaluates the LLM’s clinical diagnostic capabilities in real cases and its analysis and interpretation abilities. Clinical Diagnosis includes 6 sub-tasks, each designed according to the standard diagnostic process recommended by the Chinese official record [22] and using the chain-of-thought pattern. The 6 sub-tasks are defined as follows.
Preliminary Diagnosis (PD): The model generates a list of possible diseases $\mathcal{P}$ based on $\mathcal{CC}$, $\mathcal{MH}$, and $\mathcal{PE}$, i.e., $\mathcal{P}=\text{PD}(\mathcal{CC,MH,PE})$. This task primarily evaluates the model’s information synthesis and reasoning abilities.
Diagnostic Basis (DB): The model needs to provide supportive medical evidence $\mathcal{B}$ for each possible disease in $\mathcal{P}$, i.e., $\mathcal{B}=\text{DB}(\mathcal{E,P})$. This task mainly evaluates the model’s evidence extraction ability and helps mitigate hallucinations.
Differential Diagnosis (DD): After considering $\mathcal{B}$, the model should perform a differential analysis $\mathcal{D}$ to exclude other diseases that have different causes but similar manifestations, i.e., $\mathcal{D}=\text{DD}(\mathcal{E,P,B})$. This task primarily evaluates the model’s abilities in analytical comparison and decision-making.
Final Diagnosis (FD): Integrating all information, the model needs to determine the final diagnosis $\mathcal{F}$, i.e., $\mathcal{F}=\text{FD}(\mathcal{E,P,B,D})$. This task mainly evaluates the model’s comprehensive judgment ability and the accuracy of clinical diagnosis.
Principle of Treatment (PT): The model determines the treatment principles and guidelines $\mathcal{G}$ for $\mathcal{F}$, i.e., $\mathcal{G}=\text{PT}(\mathcal{F})$. This task primarily evaluates the model’s memorization of medical standard operating procedures and knowledge points.
Treatment Plan (TP): The model formulates specific treatment steps and plans $\mathcal{T}$ for the given case $\mathcal{E}$, including medication, surgical intervention, and physical therapy, i.e., $\mathcal{T}=\text{TP}(\mathcal{E,P,B,D,F,G})$. This task mainly evaluates the model’s abilities in knowledge application and strategic planning.